# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v1

@@ 3.1. Problem Setting
We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time periods
@@
We define the set of user groups as $\mathcal{U}$, set of machine types $\mathcal{M}$ and set of service quality tiers $\mathcal{Q}$. Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group. Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$. The deployment $D^{i}$ must be sufficient to meet the number of incoming requests which are allocated to different service quality tiers ($A^{i}$). The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval. Note that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines. Table 1 shows an overview of all input and decision variables.
@@
$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\left(k_{m,q}\right)$ that machine $m$ can serve at quality $q$.
$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\left(p^{i}_{m,q}\right)$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
@@
$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\left(d^{i}_{m,q}\right)$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\left(a^{i}_{u,q}\right)$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
@@ 3.2. Quality of Responses
We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by different service quality tiers. In our analysis, we focus on a single user group, such as best-effort users, whose requests can be fulfilled by two distinct tiers that differ in both response quality and energy demand. In this scenario, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2. When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2.
In the following, we introduce several metrics to establish a more general definition of QoR.
Service level indicators
Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), we define the service level indicator $\text{SLI}(\alpha,\omega)\in[0,1]^{\left|U\right|\times\left|Q\right|}$. It describes the observed distribution of requests across different service tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
@@
Service level objectives
To describe the minimum QoR requirements that a service must satisfy, users of this framework can define a service-level objective $\text{SLO}\in[0,1]^{\left|U\right|\times\left|Q\right|}$. The SLO indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$.
Similarly, we define a service level upper-bound $\text{SLU}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality. For example, LLM service providers often allow non-paying users a limited number of requests to access their highest-quality models, but do not intend to serve all requests at this quality tier.
[…]
$\text{SLO}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{SLU}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
describes that requests by premium users must always be served by Tier 2, while 60-100 % of requests best-effort users are served by Tier 1 and up to 40 % by Tier 2.
As all requests for each user group have to be served by some quality
(3) $\sum_{q\in\mathcal{Q}}\text{slo}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\text{slu}_{u,q}=1\quad\forall u\in\mathcal{U}.$
General definition of QoR
We define the service’s QoR within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between SLI (Equation 2), SLO and SLU:
(4) $\text{QoR}(\alpha,\omega)=1-\min_{u\in\mathcal{U},q\in\mathcal{Q}}\left(\frac{\left|\text{sli}_{u,q}(\alpha,\omega)-\text{slo}_{u,q}\right|}{\left|\text{slu}_{u,q}-\text{slo}_{u,q}\right|}\right),$
meaning that the largest deviation from SLU will determine the final QoR. Following the above example, an
$\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$
would result in a $\text{QoR}(\alpha,\omega)=0.75$.
Validity periods
As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods $[\alpha,\alpha+1,...,\omega]$. These periods can either be disjoint (e.g., quarters, months, weeks) or—more commonly—overlapping. For the remainder of this paper, we assume that QoR is assessed over rolling windows of length $\gamma$ (integer multiple of adjustment interval $\Delta$). Staying with the above example, when the validity period is set to $\gamma=24~{}\text{h}$ and we want to guarantee a specific $\text{QoR}_{\text{target}}=0.5$, we have to ensure that within each 24 hour window at least 50 % of all requests are served by Tier 2. Formally, $0.5\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$.
Figure 4 illustrates how varying the length of validity periods influences the distribution of hourly QoR over time. Longer validity periods provide greater flexibility in balancing service quality in response to fluctuations in carbon intensity, but can lead to prolonged periods of high or low hourly QoR.
@@ 3.3. Power Model
To determine the operational carbon emissions from powering servers, we first define the power usage $p^{i}_{m,q}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$. We describe two power models from the perspective of a data center operator and a cloud tenant that report their Scope 2 and 3 emissions, respectively.
Data center operator perspective
@@
Cloud tenant perspective
Service providers that rent their infrastructure from cloud providers, indirectly report the cloud data center’s power usage as part of their Scope 3 emissions. Although cloud providers recently started reporting emissions caused by renting VMs, their methods are inconsistent, lack transparency, and are subject to change. Until providers (are required to) adopt a standardized, publicly available methodology, we are left with estimating this factor. We assume that current methodologies attribute a constant factor $P_{m}^{\text{ estimate}}$ to each machine $m$ which does not change over time or depend on the executed workload or utilization of the machine:
(7) $p^{i}_{m,q}=p_{m}^{\text{estimate}}$
@@ 3.4. Carbon Emission Model
We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ in respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied carbon $C_{m}^{\text{ emb}}$ (in gCO2)
@@
Depending on methodology, embodied emissions might not always be reported with respect to the usage of infrastructure, in which case $C_{m}^{\text{emb}}=0$.
@@ 3.5. Optimization Problem
We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies the $\text{QoR}_{\text{target}}$ across validity periods.
@@
  • Sufficient resources (11) ensures that the provided machines have sufficient capacity to serve the requests allocated to each quality tier.
@@
Note, that this model assumes perfect knowledge of the future (in particular perfect forecasts for all $C^{i}$ and $R^{i}$) and can therefore not be solved in practice. However, we use it as an upper bound for assessing potential and as the foundation of the online proposed in Section 4.
Complexity
From a cloud tenant perspective (Scope 3), the optimization problem is an NP-hard mixed-integer linear program (MILP), as the objective and all constraints are linear, but involve the discrete decision variables $D^{i}$.
From the perspective of data center operators, the Scope 2 power model from Equation (5) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$. This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve. However, during optimization, we can relax Equation (5) to the same form as Equation (7), by assuming $p^{i}_{m,q}=p_{m,q}^{\text{max}}$, as long as $n\leq 1$, i.e. the power model is concave. If the power law relationship is strictly convex, i.e. $n>1$, it becomes more efficient to deploy multiple machines at lower utilization levels $util^{i}_{q}$ rather than fully utilizing fewer machines. However, this case cannot be observed in cloud data centers, where a main objective for energy efficiency is VM consolidation, to minimize the number of active machines (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007). Hence, for $n\leq 1$, the power models for Scope 2 and Scope 3 reporting are equivalent during optimization, as they result in the same decisions for optimal $A^{i}$ and $D^{i}$.
Lastly, when values in $D^{i}$ are sufficiently big (i.e. in large-scale scenarios with many active machines), the optimal result can be approximated by relaxing $D^{i}$ to continuous variables. This turns the problem into a simple linear program that is solvable in polynomial time.
@@ 4. Online Approach
For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$ as time progresses. Its performance is limited by two factors:
@@
    the errors of request and carbon intensity forecasts
@@
    the required approximations for solving the NP-hard optimization problem
Yet, our experiments show that even under realistic forecasts and strict time limits for solving the MILP, we can achieve close-to-optimal emission savings (Section 5.4).
4.1. Forecasting
We denote forecasts for request volume and carbon intensity as $\widehat{R}^{i}$ and $\widehat{C}^{i}$, respectively. There are a variety of different approaches commonly for long-term forecasts with multiple seasonalities, like Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011). As the electricity mix of regions does usually not change abruptly, long-term forecasts for ACI can be estimated reasonably well by fitting daily, weekly, and yearly seasonalities. However, especially in regions that heavily rely on weather-dependent renewable power production, additionally utilizing short-term forecasts (typically up to 4 days) provides significantly better accuracy (Maji et al., 2023; Li et al., 2024c). For example, CarbonCast (Maji et al., 2023) reports mean absolute percentage errors (MAPE) of only 3-10 % for 24-hour forecasts.
@@ 4.2. Multi-Horizon Optimization
Based on the optimization problem introduced in Section 3.5, we introduce a multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$. The approach is described in Algorithm 1.
@@
7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}}\sum_{i=\alpha}^{I-1}E^{i}$
@@
As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality. To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
@@
    A short-term optimization, executed every interval, solves the problem within a fixed horizon [^1] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization). In Line 7 we assume $\gamma$ as the horizon.. As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma/\Delta}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly. If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies the Sufficient resources (11) constraint.
@@
4.3. Dynamic QoR Adaptation for Best-Effort Users
The optimization problem introduced in Section 3.5 minimizes emissions under a given $\text{QoR}_{\text{target}}$. However, especially for best-effort users, another relevant question is: What are the expected yearly emissions from this user group, and can we keep them within a predefined budget?
In Section 5.5, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected yearly emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year. Building on this, we propose a variation of the online approach introduced in Section 4, which finds an optimal $\text{QoR}_{\text{target}}$ for best-effort users to stay within a carbon budget $B$:

### v2

@@ 3.1. Problem Setting
We define an optimization time window $\left[0,T\right]$. Given an interval $\Delta$, we divide the window into $T/\Delta$ time intervals
@@
We define the set of user groups as $\mathcal{U}$, machine types $\mathcal{M}$ and service quality tiers $\mathcal{Q}$. Machine types can represent physical hardware as well as virtual machines (VMs) like cloud instances.
For each time interval, there is an associated carbon intensity $C^{i}$ and a number of requests $R^{i}$ per user group. Each machine can handle a certain number of requests by serving a certain quality tier (denoted as $K$), which corresponds to its current power usage $P^{i}$. The deployment $D^{i}$ must be sufficient to meet the number of incoming requests that are allocated to different service quality tiers ($A^{i}$) [^1] Note, that deployments can be constrained, i.e., $d^{i}_{m,q}\in[0,\dotsc,\bar{d}]$, where $\bar{d}$ represents the maximum number of available machines. Furthermore, when elements in $D^{i}$ are sufficiently large (i.e. scenarios with many active machines), an optimal result can be approximated by relaxing $D^{i}$ to continuous variables. This turns the optimization problem introduced later into a linear program that can be solved in polynomial time.. The key challenge lies is determining the optimal values for both $D^{i}$ and $A^{i}$ for each interval. Table 1 shows an overview of all input and decision variables.
@@
$P^{i}\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Power usage $\bigl{(}p^{i}_{m,q}\bigl{)}$ of machine $m$ serving quality $q$ during $\mathcal{T}_{i}$.
$K\in\mathbb{R}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Requests per timestep $\bigl{(}k_{m,q}\bigl{)}$ that machine $m$ can serve at quality $q$.
@@
$D^{i}\in\mathbb{N}^{\left|\mathcal{Q}\right|\times\left|\mathcal{M}\right|}$ Deployment $\bigl{(}d^{i}_{m,q}\bigl{)}$, i.e., the number of machines $m$ running quality $q$ during $\mathcal{T}_{i}$.
$A^{i}\in\mathbb{R}^{\left|\mathcal{U}\right|\times\left|\mathcal{Q}\right|}$ Number of requests $\bigl{(}a^{i}_{u,q}\bigl{)}$ served by quality $q$ for user group $u$ during $\mathcal{T}_{i}$.
@@ 3.2. Quality of Responses
We define $\text{QoR}\in[0,1]$ as a metric that describes the proportion of requests served by different service quality tiers. In our subsequent analysis, we concentrate on a single user group, such as best-effort users, whose requests can be served by two distinct service tiers that differ in both response quality and energy demand. In this scenario, $\text{QoR}=0$ indicates that all requests are served by the lower-quality Tier 1, while $\text{QoR}=1$ indicates that all requests are served by the higher-quality but more energy-intensive Tier 2. When $\text{QoR}=0.5$, half of the requests are served by Tier 1 and half by Tier 2.
In the following, we introduce several metrics to establish a more general definition of QoR, that holds for multiple user groups and service quality tiers.
Service level objectives.  Adopting common terminology for quantifying QoS metrics (Keller and Ludwig, 2003), users of this framework can define a lower-bound service level objective $\underline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$, where $\underline{\text{\text{slo}}}_{u,q}$ indicates the minimal proportion of requests that should be served by quality tier $q$ for user group $u$. Similarly, users can define an upper-bound $\overline{\text{SLO}}\in[0,1]^{\left|U\right|\times\left|Q\right|}$ to service quality. For example, LLM service providers often allow non-paying users a limited number of requests to access their highest-quality models, but do not intend to serve all requests at this quality tier.
[…]
$\underline{\text{SLO}}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\quad\text{and}\quad\overline{\text{SLO}}=\begin{bmatrix}0&1\\ 0.6&0.4\end{bmatrix}$
describes that premium users’ requests must always be served by Tier 2, while 60–100 % of requests by best-effort users can be served by Tier 1 and 0–40 % by Tier 2.
All requests per user group must be served by a quality tier:
(2) $\sum_{q\in\mathcal{Q}}\underline{\text{slo}}_{u,q}=1\quad\text{and}\quad\sum_{q\in\mathcal{Q}}\overline{\text{slo}}_{u,q}=1\quad\forall u\in\mathcal{U}.$
Service level indicators.  The service level indicator $\text{SLI}(\alpha,\omega)$ describes the observed distribution of requests across different quality tiers within the interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$:
@@
As the service level is constrained by the SLOs, each
(4) $\text{sli}_{u,q}(\alpha,\omega)\in\left[\min(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q}),\max(\underline{\text{slo}}_{u,q},\overline{\text{slo}}_{u,q})\right].$
General definition of QoR.  We define the service’s quality of responses $\text{QoR}(\alpha,\omega)$ within an interval $\bigcup_{i=\alpha,\dotsc,\omega}\mathcal{T}_{i}$ through the relation between $\text{SLI}(\alpha,\omega)$, $\underline{\text{SLO}}$ and $\overline{\text{SLO}}$. In particular, for the set of parameters that permit optimization, defined as $\mathcal{V}=\{(u,q)\in\mathcal{U}\times\mathcal{Q}:\overline{\text{slo}}_{u,q}\neq\underline{\text{slo}}_{u,q}\}$, we define QoR as:
(5) $\text{QoR}(\alpha,\omega)=1-\min_{(u,q)\in\mathcal{V}}\left(\frac{\big{|}\overline{\text{slo}}_{u,q}-\text{sli}_{u,q}(\alpha,\omega)\big{|}}{\big{|}\overline{\text{slo}}_{u,q}-\underline{\text{slo}}_{u,q}\big{|}}\right).$
Following the above example, an $\text{SLI}(\alpha,\omega)=\begin{bmatrix}0&1\\ 0.7&0.3\end{bmatrix}$ would result in a $\text{QoR}(\alpha,\omega)=0.75$.
Validity periods.  As is typical in the assessment of QoS metrics (Keller and Ludwig, 2003), QoR is defined over validity periods $[\alpha,\alpha+1,...,\omega]$. These periods can either be disjoint (e.g., quarters, months, weeks) or—more commonly—overlapping. For the remainder of this paper, we assume that QoR is assessed over rolling windows of length $\gamma$ (integer multiple of adjustment interval $\Delta$). Staying with the above example, when the validity period is set to $\gamma=24~{}\text{h}$ and we want to guarantee a specific $\text{QoR}_{\text{target}}=0.5$, we have to ensure that within each 24 hour window at least 50 % of all requests are served by Tier 2. Formally, $0.5\leq\min\left\{\text{QoR}(i,i+\gamma)\right\}_{i=\alpha}^{\omega}$.
Figure 3 illustrates how the validity period length influences the hourly QoR over time. Longer validity periods provide greater flexibility in balancing service quality in response to fluctuations in carbon intensity, but can lead to prolonged periods of high or low hourly QoR.
@@ 3.3. Power Model
To determine the operational carbon emissions from powering servers, we define the power usage $P^{i}=\bigl{(}p^{i}_{m,q}\bigl{)}$ of a machine $m$ running service quality $q$ during interval $\mathcal{T}_{i}$. We describe two power models from the perspective of a data center operator and a cloud tenant that report their Scope 2 and 3 emissions, respectively.
@@
Cloud tenant perspective.  Service providers that rent their infrastructure from cloud providers, indirectly report the cloud data center’s power usage as part of their Scope 3 emissions. Although cloud providers recently started reporting emissions caused by renting VMs, their methods are inconsistent, lack transparency, and are subject to change. Until providers (are required to) adopt a standardized, publicly available methodology, we are left with estimating this factor. We assume that cloud providers attribute a constant, load-independent factor $P_{m}^{\text{attr}}$ to each active machine $m$:
(8) $p^{i}_{m,q}=p_{m}^{\text{attr}}$
@@ 3.4. Carbon Emission Model
We define a service’s estimated carbon emissions $E^{i}$ during time period $\mathcal{T}_{i}$ with respect to the region’s current operational carbon intensity $C^{i}$ (in gCO2/kWh) and the machine’s embodied emissions $C_{m}^{\text{ emb}}$ (in gCO2):
@@
Depending on methodology, embodied emissions might not always be reported with respect to infrastructure usage, in which case $C_{m}^{\text{emb}}=0$.
@@ 3.5. Optimization Problem
We formalize the problem of optimizing service deployments $D^{i}$ that minimize carbon emissions, while providing an allocation $A^{i}$ of requests to different service quality tiers that satisfies $\text{QoR}_{\text{target}}$:
@@
  • Sufficient resources (12) ensures that the provided machines have sufficient capacity to serve all requests per quality tier.
  • Respect SLOs (13) ensures that all SLOs are respected.
@@
Note, that this model assumes knowledge of future carbon intensity ($C^{i}$) and requests ($R^{i}$) and can therefore not be solved in practice. However, we use it for assessing upper-bound potential and as the foundation of the online approach proposed in Section 4.
3.6. Complexity
The proposed optimization problem is an NP-hard mixed-integer linear program (MILP), as we minimize a linear objective function subject to linear constraints, which include discrete decision variables in $D^{i}$.
Equivalence of power models.  From the perspective of data center operators, the load-dependant power model from (6) introduces a non-linear relationship by multiplying values within $D^{i}$ and $A^{i}$. This would turn the MILP into an mixed-integer quadratic program, which is considerably harder to solve. However, during optimization, we can relax (6) to the same form as (8), as long as $n\leq 1$, i.e. the power model is concave.
Theorem 3.1.
We assume $p_{m,q}^{\text{max}}>p_{m}^{\text{idle}}>0$ and $n\in[0,1]$. When minimizing (10) subject to (11) – (14) from the perspective of data center operators, the utilization-dependent power model,
$p^{i}_{m,q}=p_{m}^{\text{idle}}+\bigl{(}p_{m,q}^{\text{max}}-p_{m}^{\text{idle}}\bigl{)}\cdot\bigl{(}\text{util}^{i}_{q}\bigl{)}^{n},$
can be simplified to
$p^{i}_{m,q}=p_{m,q}^{\text{max}},$
as it always yields an equivalent optimal deployment $D^{i}$.
The proof is provided in Appendix A. Intuitively, only if the power-law relationship is strictly convex ($n>1$), it can be more efficient to deploy multiple machines at lower utilization rather than fully utilizing fewer machines. However, this case cannot be observed in cloud data centers, where a main objective for energy efficiency is VM consolidation, to minimize the number of active machines (Xu et al., 2017; Gill and Buyya, 2018; Barroso and Hölzle, 2007). Hence, for $n\leq 1$, we can use a load-independent power model during optimization, as it yields the same optimal $D^{i}$.
@@ 4. Online Approach
For assessing the practicability of carbon-aware QoR adaptation under realistic conditions, we propose an online approach, which iteratively determines the optimal $A^{i}$ and $D^{i}$ for every interval $\mathcal{T}_{i}$. Its performance is limited by two factors:
@@
    the error of request and carbon intensity forecasts
@@
    the error introduced by approximating the NP-hard optimization problem
In our experiments, conducted with realistic forecasts and strict time limits for solving the MILP, we achieve emission savings of $82\pm 6$ relative to the upper-bound potential (Section 5.4).
@@ 4.1. Multi-Horizon Optimization
Based on the optimization problem introduced in Section 3.5, we propose a forecast-based multi-horizon optimization which determines the optimal deployment and load balancing for every $\mathcal{T}_{i}$.
Forecasting.  There exist various approaches for long-term load forecasting, such as Prophet (Taylor and Letham, 2017), DeepAR (Flunkert et al., 2017), or TBATS (De Livera et al., 2011). As the electricity mix of regions does not change abruptly, long-term forecasts for ACI can be usually estimated reasonably well by fitting daily, weekly, and annual seasonalities. However, especially in regions that are heavily dependent on weather-dependent renewable power production, additionally using short-term forecasts can provide significantly higher accuracy (Maji et al., 2023; Li et al., 2024c). For example, CarbonCast (Maji et al., 2023) reports mean absolute percentage errors (MAPE) of only 3-10 % for 24-hour forecasts.
Algorithm.  The approach is described in Algorithm 1. As the optimization problem is computationally hard, we approximate near-optimal solutions by either imposing time limits or terminating the optimization process once it reaches a solution within an acceptable proximity to optimality. To reduce the negative impact of these approximations on our decisions, we perform the optimization in two distinct steps with different levels of precision:
@@
7:   $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}\leftarrow\min_{(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}}\sum_{i=\alpha}^{I-1}E^{i}$
@@
    A short-term optimization, executed every interval, solves the problem within a fixed horizon [^2] We found that the validity period length $\gamma$ is a good indicator for horizon length, although it can be shorter or longer (as future values of $A^{i}$ are fixed by the long-term optimization). In Line 7 we assume $\gamma$ as the horizon.. As we only optimize over $(D^{i},A^{i})_{i=\alpha}^{\alpha+\gamma}$, this optimization involves significantly fewer variables and is more likely to find an optimal or near-optimal solution quickly. If no solution is found, we select an $A^{i}$ such that $\text{QoR}(\alpha,\omega)=1$, and determine the minimal $D^{i}$ that satisfies the Sufficient resources (12) constraint.
@@
4.2. Automatic QoR Adaptation
The optimization problem introduced in Section 3.5 minimizes emissions under a given $\text{QoR}_{\text{target}}$. However, especially for best-effort users, another relevant question is: What are the expected annual emissions from this user group, and can we keep them within a predefined budget?
In Section 5.4, we show how running multiple long-term optimizations can provide service providers with valuable insights into expected annual emissions, allowing for adjustments to $\text{QoR}_{\text{target}}$ throughout the year. Building on this, we propose a variation of the online approach introduced in Section 4, which finds an optimal $\text{QoR}_{\text{target}}$ for best-effort users to stay within a carbon budget $B$: